OpenStudy (anonymous):

If i=square root of -1, where i be an imaginary number. Show that i=0 or i>0 or i<0? Plz can you all solve for me?

OpenStudy (ash2326):

i is a complex no., 0 is a real no. We can't compare them directly

OpenStudy (anonymous):

yeah.

OpenStudy (ash2326):

If you take the modulus of i \[|i|=\sqrt{1^2}=1\] Then \[|i|>1\]

OpenStudy (ash2326):

oops \[|i|>0\]

OpenStudy (anonymous):

and so what for i<0?

OpenStudy (ash2326):

\[i^2=-1\] so \[i^2<0\]

OpenStudy (anonymous):

Thank you so much! ash2326

OpenStudy (ash2326):

:)